# Linear Phase Balancing Scheme using Voltage Unbalance Sensitivities in Multi-phase Power Distribution Grids

### v1

@@ III-C2 LVUR sensitivities (LVURS)
can be computed similar to PVURS. The key difference here is that we need the sensitivity of line voltages magnitudes. The sensitivity of line voltages for phase $\phi,\psi,\psi^{\prime}\in\Phi$ as $v_{i\psi\psi^{\prime}}$ (complex) can be computed as $\frac{\partial v_{i\psi\psi^{\prime}}}{\partial x_{j\phi}}=\frac{\partial v_{i\psi}}{\partial x_{j\phi}}-\frac{\partial v_{i\psi^{\prime}}}{\partial x_{j\phi}}$ and the sensitivity of line voltage magnitudes can be computed as
@@
$&\frac{\partial(\Xi_{i}^{\text{LVUR}})}{\partial x_{j\phi}}=\frac{1}{\big{(}\sum_{\psi\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\big{)}^{2}}\Bigg{\{}\Big{(}\sum_{\psi\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\Big{)}\times\\ &\texttt{sgn}{(2|v_{iab}|-|v_{ibc}|-|v_{ica}|)}\times\Bigg{(}2\frac{\partial|v_{iab}|}{\partial x_{j\phi}}-\frac{\partial|v_{ibc}|}{\partial x_{j\phi}}-\\ &\frac{\partial|v_{ica}|}{\partial x_{j\phi}}\Bigg{)}-\big{|}2|v_{iab}|-|v_{ibc}|-|v_{ica}|\big{|}\times\Bigg{(}\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}\frac{\partial{|v_{i\psi\psi^{\prime}}|}}{\partial x_{j\phi}}\Bigg{)}\Bigg{\}}.\end{aligned}$ (15)
@@
$&\frac{\partial(\Xi_{i}^{\text{LVUR}})}{\partial x_{j\phi}}=\frac{1}{\big{(}\sum_{\psi\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\big{)}^{2}}\Bigg{\{}\Big{(}\sum_{\psi\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\Big{)}\times\\ &\texttt{sgn}{(2|v_{ibc}|-|v_{iab}|-|v_{ica}|)}\times\Bigg{(}2\frac{\partial|v_{ibc}|}{\partial x_{j\phi}}-\frac{\partial|v_{iab}|}{\partial x_{j\phi}}-\\ &\frac{\partial|v_{ica}|}{\partial x_{j\phi}}\Bigg{)}-\big{|}2|v_{ibc}|-|v_{iab}|-|v_{ica}|\big{|}\times\Bigg{(}\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}\frac{\partial{|v_{i\psi\psi^{\prime}}|}}{\partial x_{j\phi}}\Bigg{)}\Bigg{\}}.\end{aligned}$ (16)
@@
$&\frac{\partial(\Xi_{i}^{\text{LVUR}})}{\partial x_{j\phi}}=\frac{1}{\big{(}\sum_{\psi\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\big{)}^{2}}\Bigg{\{}\Big{(}\sum_{\psi\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\Big{)}\times\\ &\texttt{sgn}{(2|v_{ica}|-|v_{ibc}|-|v_{iab}|)}\times\Bigg{(}2\frac{\partial|v_{ica}|}{\partial x_{j\phi}}-\frac{\partial|v_{ibc}|}{\partial x_{j\phi}}-\\ &\frac{\partial|v_{iab}|}{\partial x_{j\phi}}\Bigg{)}-\big{|}2|v_{ica}|-|v_{ibc}|-|v_{iab}|\big{|}\times\Bigg{(}\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}\frac{\partial{|v_{i\psi\psi^{\prime}}|}}{\partial x_{j\phi}}\Bigg{)}\Bigg{\}}.\end{aligned}$ (17)
In the following, the sensitivity of the unbalance metrics are used for voltage balancing.
@@ IV-A Linearization using sensitivity coefficients
Using the sensitivities defined in the previous section, generic voltage unbalance metrics $(\Xi_{i}^{\theta})$ where $\theta=\{\text{VUF},\text{PVUR},\text{LVUR}\}$ can be expressed as a linear expression for each node $i\in\mathcal{N}$ using the first order Taylor’s approximation. Let’s assume that the control variables i.e., the power injection change by $\mathbf{dx}$, then the voltage unbalance metrics can be computed as
@@
In similar way, the voltage magnitudes can be linearized using the voltage sensitivity coefficients [26] as
$|v_{i\phi}(\mathbf{x+dx})|$ $\approx|v_{i\phi}({\mathbf{x}})|+\sum_{j}\sum_{\phi}\frac{\partial|v_{i\phi}(\mathbf{x})|}{\partial x_{j\phi}}(dx_{j\phi}).$ (19)
@@ IV-B Phase balancing optimization problem
In a three-phase unbalanced system, phase balancing mechanism usually refers the balancing of the voltage magnitudes, i.e., the three-phases become nearly balanced such that the overall unbalancedness comply with the limits defined by IEEE [2], IEC [3] or NEMA [4] as defined in Sec. III. In this specific case, we demonstrate the phase balancing application using active/reactive power control. Let consider that at each phase ($\phi$) of each node ($j$), there is flexibility to modify the injection by $dx_{j\phi}$ which is bounded by a factor $\beta\in[0,0.05]$ of the nominal base load per phase per node which means that the active or reactive power flexibility is limited by 5%.
@@
where $\overline{\Xi}^{\theta}$ is the upper unbalance metric defined by standards [2, 3, 4] which in most cases is 0.02 per unit as also described in Sec. III. The symbols $\underline{v}$ and $\overline{v}$ are lower and upper bounds on the voltage magnitudes which are set as 0.95 and 1.05 per unit, respectively. Eq. (20d) are the limits on the controllable power flexibility per phase per node, where $\hat{\mathbf{x}}$ refers to the nominal injection. Eq. (20e) ensures that the sum of the new power injections matches the net demand per node in the original system, which means that net sum of flexible deviations are zero.
As it can be seen, this phase balancing formulation is linear program (LP), thanks to the linearized expression in (18) and (19). The proposed formulation will be validated in the following section on an IEEE benchmark network.
@@ V Numerical Validation
(a) $\partial(\Xi_{\text{VUF}})/\partial p_{a}$
(b) $\partial(\Xi_{\text{VUF}})/\partial p_{b}$
(c) $\partial(\Xi_{\text{VUF}})/\partial q_{a}$
(d) $\partial(\Xi_{\text{VUF}})/\partial q_{b}$
(e) $\partial(\Xi_{\text{PVUR}})/\partial p_{a}$
(f) $\partial(\Xi_{\text{PVUR}})/\partial p_{b}$
(g) $\partial(\Xi_{\text{PVUR}})/\partial q_{a}$
(h) $\partial(\Xi_{\text{PVUR}})/\partial q_{b}$
(i) $\partial(\Xi_{\text{LVUR}})/\partial p_{a}$
(j) $\partial(\Xi_{\text{LVUR}})/\partial p_{b}$
(k) $\partial(\Xi_{\text{LVUR}})/\partial q_{a}$
(l) $\partial(\Xi_{\text{LVUR}})/\partial q_{b}$
@@
(a) $\partial(\Xi_{\text{VUF}})/\partial p$
(b) $\partial(\Xi_{\text{VUF}})/\partial q$
Figure 3: VUF sensitivity with varying $p,q$.
(a) $\partial(\Xi_{\text{PVUR}})/\partial p$
(b) $\partial(\Xi_{\text{PVUR}})/\partial q$
Figure 4: PVUR sensitivity with varying $p,q$.
(a) $\partial(\Xi_{\text{LVUR}})/\partial P$
(b) $\partial(\Xi_{\text{LVUR}})/\partial Q$
Figure 5: LVUR sensitivity with varying $p,q$.
In the following, first, we present the results for sensitivity coefficient assuming a nominal demand (Sec. V-A), then we vary the injections for a particular phase and node to observe its effect on the sensitivity coefficient (Sec. V-B), then finally, we present the result of phase balancing scheme (Sec. V-C).
@@ V-A Sensitivity coefficients per phase per node
Figure 2 shows VUFS, PVURS and LVURS for IEEE-34 case in the form of heatmap plot. These sensitivities are computed for each node with respect to the active and reactive power injections of different nodes and phases in Figs 2(a-l). Here $p_{a}$, $p_{b}$, $q_{a}$ and $q_{b}$ denote active powers of phases $a$ and $b$ and reactive power at phase $a$ and $b$, respectively. As it can be seen, these sensitivities vary significantly along the nodes and phases and differ for different unbalance metrics. From the plots, it can be observed that the VUFS and PVURS vary over a wide range of values, whereas the LVURS remain fairly within a small range, which may suggest that LVUR is equally sensitive to the power injections irrespective of their location in the network, wheres as VUF and PVUR are quite sensitive with respect to the locations of the injections.
@@ V-B Sensitivity with change in active and reactive injections
Figures 3, 4 and 5 report the evolution of the VUF, PVUR and LVUR sensitivities with change in the injections of one particular node and phase. We present the sensitivities of nodes 802, 816, 828, 848 and 858 with respect to the active and reactive injections at node 862, phase $c$. The active and reactive powers are varied together by multiplication factor to the nominal injection at that bus from -1.5 to 1.5 with intervals of 0.1. As seen from the plot, the sensitivity of VUF with respect to active and reactive powers are smooth, whereas the ones for PVUR and LVUR are discontinuous. This is because, the PVUR and LVUR sensitivities derived in Sec.III-B2 and III-C2 were not smooth. The plot also shows that these sensitivities are changing rapidly with the change in the injections, which tells that the linearization of the VUF, PVUR and LVUR are only valid in close vicinity of the operating point.
@@
V-C Phase balancing application
This section illustrates how the sensitivity-based linear phase balancing formulation in Sec. IV-B can improve unbalance metrics by reallocating power injections per phase per node. The objective is to minimize a specific unbalance metric of interest (i.e., VUF, PVUR or LVUR) with respect the voltage constraints and the bounds of the unbalance metric. Figure. 6 shows the flow diagram for the phase balancing scheme. As it is shown, using the known operating point (either from the measurements followed or by a state estimation process), first the coefficients for the voltage magnitudes and the unbalance metrics are computed, then they are used in the phase balancing scheme in (20), then the setpoints are realized and the response is observed in terms of actualized voltages, which acts as the operating point for the next time step.
@@ V-C Phase balancing application
Figure 7: Voltage balancing results for $\beta=$ 0.01 and 0.02. Base case refers to no phase balancing, Optimized is refers to the results after solving phase balancing problem in (20) and Realized is the actualized value after solving AC power flow ex-post with outputs of phase balancing problem.
TABLE I: Phase balancing results
Metrics No balancing Balancing Balancing
$\beta=0.01$ $\beta=0.02$
@@ VI Conclusion
This work proposed a linear phase balancing scheme for distribution systems to tackle the inherent non-linearity of voltage unbalance metrics and power flow equations. It presented a new set of sensitivity coefficients related to the voltage unbalance definitions from IEEE, IEC, and NEMA, namely Voltage Unbalance Factors (VUF), Phase Voltage Unbalance Rates (PVUR), and Line Voltage Unbalance Rates (LVUR).
@@
Future work will explore the use of unbalance metric sensitivities for other applications such as phase swapping or phase allocation schemes, resulting in mixed integer problems.

### v2

@@ III-C2 LVUR sensitivities (LVURS)
can be computed similarly to PVURS. The key difference here is that we need the sensitivity of line voltage magnitudes. The sensitivity of line voltages for phase $\phi,\psi,\psi^{\prime}\in\Phi$ as $v_{i\psi\psi^{\prime}}$ (complex) can be computed as $\frac{\partial v_{i\psi\psi^{\prime}}}{\partial x_{j\phi}}=\frac{\partial v_{i\psi}}{\partial x_{j\phi}}-\frac{\partial v_{i\psi^{\prime}}}{\partial x_{j\phi}}$ and the sensitivity of line voltage magnitudes can be computed as
@@
$\begin{aligned} &\frac{\partial(\Xi_{i}^{\text{LVUR}})}{\partial x_{j\phi}}=\frac{1}{\big(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\big)^{2}}\Bigg\{\Big(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\Big)\times\\ &\texttt{sgn}{(2|v_{iab}|-|v_{ibc}|-|v_{ica}|)}\times\Bigg(2\frac{\partial|v_{iab}|}{\partial x_{j\phi}}-\frac{\partial|v_{ibc}|}{\partial x_{j\phi}}-\\ &\frac{\partial|v_{ica}|}{\partial x_{j\phi}}\Bigg)-\big|2|v_{iab}|-|v_{ibc}|-|v_{ica}|\big|\times\Bigg(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}\frac{\partial{|v_{i\psi\psi^{\prime}}|}}{\partial x_{j\phi}}\Bigg)\Bigg\}.\end{aligned}$ (15)
@@
$\begin{aligned} &\frac{\partial(\Xi_{i}^{\text{LVUR}})}{\partial x_{j\phi}}=\frac{1}{\big(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\big)^{2}}\Bigg\{\Big(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\Big)\times\\ &\texttt{sgn}{(2|v_{ibc}|-|v_{iab}|-|v_{ica}|)}\times\Bigg(2\frac{\partial|v_{ibc}|}{\partial x_{j\phi}}-\frac{\partial|v_{iab}|}{\partial x_{j\phi}}-\\ &\frac{\partial|v_{ica}|}{\partial x_{j\phi}}\Bigg)-\big|2|v_{ibc}|-|v_{iab}|-|v_{ica}|\big|\times\Bigg(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}\frac{\partial{|v_{i\psi\psi^{\prime}}|}}{\partial x_{j\phi}}\Bigg)\Bigg\}.\end{aligned}$ (16)
@@
$\begin{aligned} &\frac{\partial(\Xi_{i}^{\text{LVUR}})}{\partial x_{j\phi}}=\frac{1}{\big(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\big)^{2}}\Bigg\{\Big(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}{|v_{i\psi\psi^{\prime}}|}\Big)\times\\ &\texttt{sgn}{(2|v_{ica}|-|v_{ibc}|-|v_{iab}|)}\times\Bigg(2\frac{\partial|v_{ica}|}{\partial x_{j\phi}}-\frac{\partial|v_{ibc}|}{\partial x_{j\phi}}-\\ &\frac{\partial|v_{iab}|}{\partial x_{j\phi}}\Bigg)-\big|2|v_{ica}|-|v_{ibc}|-|v_{iab}|\big|\times\Bigg(\sum_{\psi,\psi^{\prime},\psi\neq\psi^{\prime}}\frac{\partial{|v_{i\psi\psi^{\prime}}|}}{\partial x_{j\phi}}\Bigg)\Bigg\}.\end{aligned}$ (17)
In the following, the sensitivity of the unbalance metrics is used for voltage balancing.
@@ IV-A Linearization using sensitivity coefficients
In this section, we show the use of voltage unbalance sensitivities for linearized expression of voltage unbalance factors. Let denote the generic voltage unbalance metrics by $(\Xi_{i}^{\theta})$ where $\theta=\{\text{VUF},\text{PVUR},\text{LVUR}\}$; using the sensitivities defined in the previous section, they can be expressed as a linear expression for each node $i\in\mathcal{N}$ using the first order Taylor’s approximation. Assuming that the control variables, i.e., the power injection are denoted by $\mathbf{x}$, then the voltage unbalance metrics can be computed as
@@
In a similar way, the voltage magnitudes can be linearized using the voltage sensitivity coefficients [26] as
$|v_{i\psi}(\mathbf{x+dx})|$ $\approx|v_{i\psi}({\mathbf{x}})|+\sum_{j}\sum_{\phi}\frac{\partial|v_{i\psi}(\mathbf{x})|}{\partial x_{j\phi}}(dx_{j\phi}).$ (19)
@@ IV-B Phase balancing optimization problem
In a three-phase unbalanced system, phase balancing mechanism usually refers to the balancing of the voltage magnitudes, i.e., the three-phases become nearly balanced such that the overall unbalancedness complies with the limits defined by IEEE [2], IEC [3] or NEMA [4] as defined in Sec. III. In this specific case, we demonstrate the phase balancing application using active/reactive power control. Let consider that at each phase ($\phi$) of each node ($j$), there is flexibility to modify the injection by $dx_{j\phi}$ which is bounded by a factor $\beta\in[0,0.05]$ of the nominal base load per phase per node which means that the active or reactive power flexibility is limited by 5%.
@@
where $\overline{\Xi}^{\theta}$ is the upper unbalance metric defined by standards [2, 3, 4] which in most cases is 0.02 per unit as also described in Sec . III. The symbols $\underline{v}$ and $\overline{v}$ are lower and upper bounds on the voltage magnitudes, which are set as 0.95 and 1.05 per unit, respectively. Eq. (20d) are the limits on the controllable power flexibility per phase per node, where $\hat{\mathbf{x}}$ refers to the nominal injection. Eq. (20e) ensures that the sum of the new power injections matches the net demand per node in the original system, which means that the net sum of flexible deviations is zero.
As it can be seen, this phase balancing formulation is a linear program (LP), thanks to the linearized expression in (18) and (19). The proposed formulation will be validated in the following section on an IEEE benchmark network.
@@ V Numerical Validation
(a) $\partial(\Xi^{\text{VUF}})/\partial p_{a}$
(b) $\partial(\Xi^{\text{VUF}})/\partial p_{b}$
(c) $\partial(\Xi^{\text{VUF}})/\partial p_{c}$
(d) $\partial(\Xi^{\text{VUF}})/\partial q_{a}$
(e) $\partial(\Xi^{\text{VUF}})/\partial q_{b}$
(f) $\partial(\Xi^{\text{VUF}})/\partial q_{c}$
(g) $\partial(\Xi^{\text{PVUR}})/\partial p_{a}$
(h) $\partial(\Xi^{\text{PVUR}})/\partial p_{b}$
(i) $\partial(\Xi^{\text{PVUR}})/\partial p_{c}$
(j) $\partial(\Xi^{\text{PVUR}})/\partial q_{a}$
(k) $\partial(\Xi^{\text{PVUR}})/\partial q_{b}$
(l) $\partial(\Xi^{\text{PVUR}})/\partial q_{c}$
(m) $\partial(\Xi^{\text{LVUR}})/\partial p_{a}$
(n) $\partial(\Xi^{\text{LVUR}})/\partial p_{b}$
(o) $\partial(\Xi^{\text{LVUR}})/\partial p_{c}$
(p) $\partial(\Xi^{\text{LVUR}})/\partial q_{a}$
(q) $\partial(\Xi^{\text{LVUR}})/\partial q_{b}$
(r) $\partial(\Xi^{\text{LVUR}})/\partial q_{c}$
@@
In the following, first, we present the results for sensitivity coefficient assuming a nominal demand (Sec. V-A), then we vary the injections for a particular phase and node to observe its effect on the sensitivity coefficient (Sec. V-B), then finally, we present the result of phase balancing scheme (Sec. V-D).
@@ V-A Sensitivity coefficients per phase per node
Figure 2 shows VUFS, PVURS, and LVURS for the IEEE-34 case in the form of a heatmap plot. These sensitivities are computed for each node with respect to the active and reactive power injections of different nodes and phases in Figs 2(a-l). Here $p_{a}$, $p_{b}$, $p_{c}$, $q_{a}$, $q_{b}$ and $q_{c}$ denote active powers of phases $a,b$ and $c$ and reactive power at phase $a,b$ and $c$, respectively. As it can be seen, these sensitivities vary significantly along the nodes and phases and differ for different unbalance metrics. From the plots, it can be observed that the VUFS and PVURS vary over a wide range of values, whereas the LVURS remain fairly within a small range. It suggests that LVUR is equally sensitive to the power injections irrespective of their location in the network, whereas VUF and PVUR are quite sensitive with respect to the locations of the injections.
@@ V-B Sensitivity with change in active and reactive injections
Figures 3, 4 and 5 report the evolution of the VUF, PVUR, and LVUR sensitivities with change in the injections of one particular node and phase. We present the sensitivities of nodes 802, 816, 828, 848, and 858 with respect to the active and reactive injections at node 862, phase $c$. The active and reactive powers are varied together by a multiplication factor to the nominal injection at that bus from -1.5 to 1.5 with intervals of 0.1. As seen from the plot, the sensitivity of VUF with respect to active and reactive powers is smooth, whereas the ones for PVUR and LVUR are discontinuous. This is because the PVUR and LVUR sensitivities derived in Sec.III-B2 and III-C2 were not smooth. The plot also shows that these sensitivities are changing rapidly with the change in the injections, which tells that the linearization of the VUF, PVUR, and LVUR is only valid in the close vicinity of the operating point.
(a) $\partial(\Xi^{\text{VUF}})/\partial p$
(b) $\partial(\Xi^{\text{VUF}})/\partial q$
Figure 3: VUF sensitivity with varying $p,q$ at node 862, phase c.
(a) $\partial(\Xi^{\text{PVUR}})/\partial p$
(b) $\partial(\Xi^{\text{PVUR}})/\partial q$
Figure 4: PVUR sensitivity with varying $p,q$ at node 862, phase c.
(a) $\partial(\Xi^{\text{LVUR}})/\partial p$
(b) $\partial(\Xi^{\text{LVUR}})/\partial q$
Figure 5: LVUR sensitivity with varying $p,q$ at node 862, phase c.
V-C Estimating VUF, PVUR, and LVUR using Sensitivities
This section shows how the sensitivities are used to estimate the VUF, PVUR, and LVUR for another operating point $\mathbf{x}$} using a known operating point $\hat{\mathbf{x}}$ using (18). For this simulation, $\hat{\mathbf{x}}$ is used as the nominal loading (State 0) as an operating point. We show estimation for two cases: (i) underloading: new injections are 0.8 times nominal load, and (ii) overloading: new injections are 1.2 times nominal load. The results for the two cases (state 1 and 2) are compared against the true value (computed with exact power flow) and shown in the barplot in Figs. 6 and 7 for VUF, PVUR and LVUR. As can be seen, the estimated values (using (18)) closely match the true AC power flow values for both cases.
(a) Estimated VUF for a perturbed state.
(b) Estimated PVUR for a perturbed state.
(c) Estimated LVUR for a perturbed state.
Figure 6: Underloading case as state 1 (0.8 $\times$ nominal load).
(a) VUF estimation for a perturbed state.
(b) PVUR estimation for a perturbed state.
(c) LVUR estimation for a perturbed state.
Figure 7: Overloading case as state 2 (1.2 $\times$ nominal load).
V-D Phase balancing application
@@
This section illustrates how the sensitivity-based linear phase balancing formulation in Sec. IV-B can improve unbalance metrics by reallocating power injections per phase per node. The objective is to minimize a specific unbalance metric of interest (i.e., VUF, PVUR, or LVUR) with respect to the voltage constraints and the bounds of the unbalance metric. Figure 8 shows the flow diagram for the phase balancing scheme. As it is shown, using the known operating point (either from the measurements followed or by a state estimation process), first the coefficients for the voltage magnitudes and the unbalance metrics are computed, then they are used in the phase balancing scheme in (20), then the setpoints are realized and the response is observed in terms of actualized voltages, which acts as the operating point for the next time step.
@@
Figure 9: Voltage balancing results for $\beta=$ 0.01 and 0.02. Base case refers to no phase balancing, Optimized refers to the results after solving phase balancing problem in (20) and Realized is the actualized value after solving AC power flow ex-post with outputs of phase balancing problem.
TABLE I: Phase balancing results.
Metrics Without Balancing Balancing
balancing $\beta=0.01$ $\beta=0.02$
@@ VI Conclusion
This work proposed a linear phase balancing scheme for distribution systems to tackle the inherent nonlinearity of voltage unbalance metrics and power flow equations. It presented a new set of sensitivity coefficients related to the voltage unbalance definitions from IEEE, IEC, and NEMA, namely Voltage Unbalance Factors (VUF), Phase Voltage Unbalance Rates (PVUR), and Line Voltage Unbalance Rates (LVUR).
@@
Future work will explore the use of unbalance sensitivities for other applications such as phase swapping or phase allocation schemes, resulting in mixed integer problems.